OpenStudy (anonymous):

Objects A and B each have a mass of 25 kilograms. Object A has a velocity of 5.98 meters/second. Object B is stationary. They undergo a perfectly elastic collision in one dimension. What is the total kinetic energy of the system after the collision?

OpenStudy (anonymous):

It is the same as when it started, the definition of elastic collision. KE = (1/2) m v^2 = 0.5 (25)(5.98)^2 = 447 J.